OpenStudy (anonymous):

5. Multiply. (3x + 5)(3x – 5) (1 point) 9x2 – 30x – 25 9x2 + 30x + 25 9x2 – 25 9x2 + 25 6. The area of a rectangular painting is given by the trinomial x2 – 5x – 36. What are the possible dimensions of the painting? Use factoring. (1 point) (x + 9) and (x – 4) (x + 9) and (x + 4) (x – 9) and (x – 4) (x – 9) and (x + 4)

pooja195 (pooja195):

FOIL

pooja195 (pooja195):

for number 5

pooja195 (pooja195):

(3x)(3x)+(3x)(−5)+(5)(3x)+(5)(−5)

OpenStudy (anonymous):

i dont know how to do that

pooja195 (pooja195):

thats 3x times 3x

pooja195 (pooja195):

*whats

OpenStudy (anonymous):

9

pooja195 (pooja195):

Once you multiply you get this 9x2−15x+15x−25

pooja195 (pooja195):

Combine LIKE TERMS

OpenStudy (anonymous):

is it 9x2 – 30x – 25

pooja195 (pooja195):

no -15+15 = 0

pooja195 (pooja195):

9x2−25

OpenStudy (shaik0124):

5 answer 9x2 – 25

pooja195 (pooja195):

x2 – 5x – 36. What numbers multiply to get -36 and add to -5

OpenStudy (shaik0124):

the question in the form (a+b)(a-b)=a^2-b^2 similarly (3x+5)(3x-5)=(3x)^2-5^2

OpenStudy (jack1):

nice @pooja195